OpenStudy (anonymous):

Imagine a 10kg block moving with a velocity of 20m/s to the left. Calculate the kinetic energy of this block

OpenStudy (isaiah.feynman):



OpenStudy (anonymous):

With the equation Isaiah gave the m= 10kg and v^2=20 so it would look like this k = 1/2(10)(20^2) k = 2000 I hope this helped! Make sure to close the question and award medals.

OpenStudy (anonymous):

thank you both, i wish it were possible for me to understand it

OpenStudy (isaiah.feynman):

What? IT IS POSSIBLE FOR YOU TO UNDERSTAND IT.

OpenStudy (anonymous):

What are you not getting? I am here to help with any questions you have and I'm sure @Isaiah.Feynman is too. @marie123